OpenStudy (anonymous):

1. 48 2. Imagine a pumpkin patch is a square garden, measuring the number of square feet you got in answer #1: Since your answer to #1 isn’t a perfect square, please round to the nearest perfect square and find the square root of that number. 3. Back to the pumpkin patch. You decide to put a path running diagonally across your patch. How long is your path? Hint: Think right triangle, think Pythagorean Theorem…

OpenStudy (anonymous):

7^2= 49 so 7*7

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

if the sides are 7 each then pythagorean theorem means 49+49=c^2 c=\[\sqrt{98}\]

OpenStudy (anonymous):

the perfect roots are 1 4 9 16 25 36 49 64.... and so on. it says to round and does not say up or down so closest would be 49